OpenStudy (anonymous):

if X and Y are positive integers, X>30 and X-Y>25,what is the LEAST possible value of X+Y???

OpenStudy (anonymous):

32

OpenStudy (anonymous):

x + y x>30 lets say x =31 x-y>25 y<6 y=1 x+y=32

OpenStudy (anonymous):

those hits just keep on coming

OpenStudy (anonymous):

why can't x by 30.001?

OpenStudy (anonymous):

An integer is a whole number.

OpenStudy (amistre64):

integers

OpenStudy (anonymous):

it is integer right

OpenStudy (anonymous):

oh yes, you are right!

OpenStudy (anonymous):

thanks...